OpenStudy (anonymous):

How many solutions can be found for the equation 3y + 5 − 2y = 11? Zero One Two Infinitely many

OpenStudy (anonymous):

@blm1 @Mehek14 @

OpenStudy (texaschic101):

3y + 5 - 2y = 11 y = 11 - 5 y = 6 one solution

OpenStudy (anonymous):

what does that mean

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

so its b